# Adversarial Coevolutionary Illumination with Generational Adversarial MAP-Elites

### v1

@@ Conclusion
We present a new algorithm, GAME, that illuminates adversarial problems using coevolution. In combination with a vision embedding model, GAME implements a behavior space requiring only videos of the behavior instead of handcrafted behavior descriptors. Results on an adversarial multi-agent game confirm GAME’s effectiveness in illuminating the adversarial space. The study also reveals interesting insights: bootstrapping reduces the reachable search space, thereby limiting open-endedness; and not pruning neutral mutations allows for preserving necessary stepping stones that lead to discovering the best-performing solutions. Future work will aim to demonstrate that GAME is a generalist algorithm applicable to various adversarial problems, not limited to symmetrical problems or behavior trees.

### v3

@@ Conclusion
We present GAME, a coevolutionary QD algorithm for illuminating adversarial problems. Combined with a vision embedding model as a domain-agnostic behavior space, GAME requires only videos instead of handcrafted behavior descriptors. We demonstrate GAME’s ability to illuminate three different adversarial problems and validate all its components through ablation studies in a multi-agent battle game. We show its generality by using the same code (GAME + VEM) when evolving both behavior trees for a multi-agent battle game and soft-robot morphologies in a wrestling environment. The deck building illumination also demonstrates its strength in fostering quality, while revealing limitations in covering small behavior spaces due to its multi-task formulation. This work is limited by the non-open-ended nature of the different search spaces, which prevents GAME from generating open-ended illumination. Future work should focus on applying GAME to open-ended search spaces, paving the way toward a better understanding of the emergence of open-ended adversarial coevolution.